# Misaligned incentives in the post-2024 UEFA Champions League: A managerial failure

### v5

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification for the post-2024 UEFA Champions League is not always based on sporting merit as we show: a match with misaligned incentives has barely been avoided in the 2023/24 German Bundesliga. Two straightforward solutions are recommended to safeguard the integrity of the competition. Fairness can be easily guaranteed by reversing the order of two articles in the Champions League regulation. UEFA repeatedly commits the same mistake in allocating vacant slots, which seems to indicate a clear managerial failure, especially because the solution is widely known. Thus, governing bodies in sports are called to work more closely together with the scientific community.
@@ 1 Introduction
Ensuring the fairness and integrity of sports is one of the key responsibilities of tournament organisers (Kendall and Lenten,, 2017; Devriesere et al.,, 2024). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ):
UEFA has clearly shown that we are fully committed to respecting the fundamental values of sport and to defending the key principle of open competitions, with qualification based on sporting merit, fully in line with the values and solidarity-based European sports model.
However, is the UEFA Champions League qualification based on sporting merit under any circumstances? Is there any issue in its rules that threatens the fairness of this prestigious tournament and can potentially jeopardise competition?
The current paper uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not avoid an unfair game in a domestic league. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots are assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig.
The problem is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the root of the problem resides in the allocation of vacant slots, which led to some instances of incentive incompatibility in the past, even though its solution has been known since March 2013 when the first version of Dagaev and Sonin, (2018) became available. Making the same mistake repeatedly seems to indicate a clear managerial failure.
@@ 2 Related literature
In the past, UEFA club competitions violated strategy-proofness mainly due to the unfair allocation of vacant slots. Vacant slots occur if teams can qualify for a tournament in more than one way, thus, the organiser needs to specify what happens if at least two conditions of qualification are satisfied. A straightforward example is the UEFA Europa League, where teams can enter via both their domestic championship and their domestic cup. According to Dagaev and Sonin, (2018), these tournament systems consisting of multiple round-robin and knockout tournaments are incentive incompatible if (1) there are at least two round-robin tournaments with at least one prize; (2) there is one round-robin tournament and at least one knockout tournament with at least one prize, and the vacant slots are not always filled through the round-robin tournament.
This theoretical finding might not seem especially interesting for sports policy. However, the misaligned allocation rule has been responsible for the incentive incompatibility of
@@
Fortunately, Section 4 will reveal that the third problem has been solved by the new seeding rules effective from the 2024/25 season. On the other hand, the reform has also created potentially misaligned incentives in the UEFA Champions League qualification as will be seen in Section 3.
The introduction of the UEFA Nations League has also allowed for games where a team has an incentive to lose (Csató,, 2022; Haugen and Krumer,, 2021). Even though Scelles et al., (2024) identify no such games between 2018 and 2021, the authors call for more research around optimal competition formats that provide appropriate incentives for all teams in all games. Our paper is one of these follow-up studies.
@@ 3 A real-world example
Currently, the four highest-ranked associations—including Germany—are represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders automatically qualify for the league phase. The possible vacancy is filled according to UEFA, 2024b (, Article 3.04):
@@
1 Bayer Leverkusen 27 6 0 87 23 +++64 87
2 Bayern München 23 3 7 92 41 +++51 72
3 VfB Stuttgart 22 4 7 74 39 +++35 70
4 RB Leipzig 19 7 7 75 37 +++38 64
5 Borussia Dortmund 18 9 6 66 43 +++23 63
6 Eintracht Frankfurt 11 13 9 49 48 +++1 46
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
4 Discussion and policy implications
@@
A recent literature survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper has examined the new format of the UEFA Champions League from this perspective. The rules have clearly improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
On the other hand, strategy-proofness is violated in the new format according to Section 3. What is the underlying reason and how can incentive compatibility be restored?
UEFA, 2024b (, Article 3.04) has changed compared to the previous season of 2023/24. However, under both rules, no club could have an incentive to create a vacancy in order to obtain the released slot since either the highest-ranked not already qualified club from the association of the titleholder or a club from a different association benefits from the vacancy (see the cumbersome discussion of UEFA, 2024b (, Article 3.04b)). Consequently, Article 3.04 is not responsible for incentive incompatibility on its own.
@@
On the other hand, the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA, 2024b, , Article 3.07). Thus, a team playing in the same round-robin tournament (domestic championship) as one of the (potential) titleholders can be interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder—a clear example of a tanking opportunity as we have seen above. Such misaligned incentives can arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
We offer the following remedies:
@@
Both reforms satisfy incentive compatibility. For Rule change A, the argument is provided above. For Rule change B, strategy-proofness is implied by Dagaev and Sonin, (2018, Proposition 6).
@@
Since Rule change A is closer to the original intention of UEFA, this would be our first recommendation. Rule change B favours the domestic championship of the titleholder, hence, it is expected to lead to an even higher concentration of the clubs playing in the UEFA Champions League by their associations because the highest-ranked countries are more likely to win both the UEFA Champions League and the UEFA Europa League.
If UEFA is not much worried about manipulation in a domestic league, and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. Following the previous literature, the current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, (2022), even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig has been played in the last round. The set of top teams can be relatively accurately estimated at the beginning of the season, hence, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule may lead to a tanking opportunity at the end of the season with a (much) lower probability. However, even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.

### v6

@@ Abstract
UEFA declares that it is fully committed to respecting the fundamental values of sport. However, the qualification for the post-2024 UEFA Champions League is not always based on sporting merit as we show. In particular, a match with misaligned incentives has barely been avoided in the 2023/24 German Bundesliga. Our mathematical model reveals that fairness can be guaranteed by filling all vacant slots via the domestic leagues and retaining the association constraint in the league phase draw. UEFA repeatedly commits the same mistake in treating vacancies, which indicates a clear managerial failure. Therefore, governing bodies in sports are called to work more closely together with the scientific community.
@@ 1 Introduction
Ensuring the fairness and integrity of sports is a crucial responsibility of tournament organisers (Kendall and Lenten,, 2017; Devriesere et al.,, 2024). Naturally, the Union of European Football Associations (UEFA) is also committed to these values, at least in declarations and media releases. For example, UEFA president Aleksander Čeferin has commented on the novel format of the UEFA Champions League in 2024 with the following words (UEFA, 2024a, ): “UEFA has clearly shown that we are fully committed to respecting the fundamental values of sport and to defending the key principle of open competitions, with qualification based on sporting merit, fully in line with the values and solidarity-based European sports model.
However, is the UEFA Champions League qualification based on sporting merit under all circumstances? Is there any issue in its rules that threatens the fairness of the tournament and can potentially jeopardise competition?
The current paper uncovers that, while UEFA may be committed to sporting merit in principles, their regulations do not exclude an unfair game. In particular, although the recent fundamental reform of the UEFA Champions League has restored the integrity of the seeding system, the way in which the novel European performance spots is assigned allows for a situation when a team is strictly better off by losing. These misaligned incentives almost occurred in the last round of the 2023/24 German Bundesliga in the match Eintracht Frankfurt vs. RB Leipzig. A mathematical model is presented to understand the issue and compare different allocation rules of vacant slots. Two possible solutions are offered to decrease the probability of such a problem or even entirely avoid it. It is also shown that the association constraint in the league phase draw has a role in improving fairness.
The novel case of incentive incompatibility is especially worrying because “the objective for a sport organisation should be to eliminate any game that raises potential unfairness ex-ante(Scelles et al.,, 2024, p. 535). Furthermore, the root of the problem resides in the allocation of vacant slots, which led to some instances of incentive incompatibility in the past. Making the same mistake repeatedly seems to indicate a clear managerial failure.
@@ 2 Related literature
In the past, UEFA club competitions violated strategy-proofness mainly due to the unfair allocation of vacant slots. Vacant slots occur if teams can qualify for a tournament in more than one way, thus, the organiser needs to specify what happens if two conditions of qualification are simultaneously satisfied. A straightforward example is the UEFA Europa League, where teams can enter via both their domestic championship and domestic cup. According to Dagaev and Sonin, (2018), these tournament systems consisting of multiple round-robin and knockout tournaments are incentive incompatible if (1) there are at least two round-robin tournaments with at least one prize; (2) there is one round-robin tournament and at least one knockout tournament with at least one prize, and the vacant slots are not always filled through the round-robin tournament.
This theoretical result might not seem especially interesting for sports policy. However, the misaligned allocation rule has been responsible for the incentive incompatibility of
@@
Fortunately, the third problem has been solved by the new seeding rules effective from the 2024/25 season, see Section 5. On the other hand, the reform has also created potentially misaligned incentives in the UEFA Champions League qualification as will be shown in Section 3.
The introduction of the UEFA Nations League has also allowed for matches where a team has an incentive to lose (Csató,, 2022; Haugen and Krumer,, 2021). Even though Scelles et al., (2024) identify no such games between 2018 and 2021, the authors call for more research around optimal competition formats that provide appropriate incentives for all teams in all games. Our paper can be considered as one of these follow-up studies.
@@ 3 A real-world example
The four highest-ranked associations—including Germany—are currently represented by four clubs in the UEFA Champions League. In addition, both the UEFA Champions League and UEFA Europa League titleholders automatically qualify for the league phase. The possible vacancy is filled according to UEFA, 2024b (, Article 3.04):
@@
1 Bayer Leverkusen 27 6 0 87 23 $+$64 87
2 Bayern München 23 3 7 92 41 $+$51 72
3 VfB Stuttgart 22 4 7 74 39 $+$35 70
4 RB Leipzig 19 7 7 75 37 $+$38 64
5 Borussia Dortmund 18 9 6 66 43 $+$23 63
6 Eintracht Frankfurt 11 13 9 49 48 $+$1 46
@@ 7 SC Freiburg 11 9 13 44 56 --12 42
4 A mathematical model
A tournament $(X,Y)$ is a tuple of a set of results $X$ and a ranking method $Y$ that assigns a linear order of the teams to any set of results $X$. The number of teams in tournament $(X,Y)$ is denoted by $\lvert X\rvert$. For each team $1\leq i\leq\lvert X\rvert$, $Y_{i}(X)$ is its rank under a given set of results $X$.
Let $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ be a tournament system consisting of three tournaments $(S,P)$, $(T,Q)$, $(U,R)$. The winner of tournament $(S,P)$ is denoted by $p$, that is, $1\leq p\leq\lvert S\rvert$ is the (unique) team for which $P_{p}(S)=1$ holds. This team $p$ also plays in tournament $(T,Q)$ but not in tournament $(U,R)$. The winner of tournament $(U,R)$ is denoted by $r$: $1\leq r\leq\lvert U\rvert$ is the (unique) team for which $R_{r}(U)=1$ holds.
This model can represent the qualification of the UEFA Champions League as follows. $(S,P)$ is the Champions League. $(T,Q)$ is the domestic league of the Champions League winner, denoted by $p$. $(U,R)$ is the “competition” played by all teams qualified for the champions path in the qualification of the Champions League such that the ranking method $R$ is based on the club coefficients of these teams.
The allocation rule $\mathcal{A}(\mathbf{W})$ chooses the winner of tournament $(S,P)$ and the top $q$ teams of tournament $(T,Q)$. However, a vacancy may be created if team $p$ is ranked among the top $q$ teams in tournament $(T,Q)$.
We consider three allocation rules that differ in the assignment of the vacant slot caused by team $p$. As discussed in Section 3, the current rule of the UEFA Champions League is:
$\mathcal{A}(\mathbf{W})=\begin{cases}i:i=p\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\geq q+1\\ i:i=r\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\leq q-1\\ i:Q_{i}(T)\leq q+1&\text{if }Q_{p}(T)=q.\end{cases}$
Note that the winner $r$ of tournament $(U,R)$ receives the vacant slot only if the titleholder team $p$ is ranked among the top $q-1$ teams in tournament $(T,Q)$ because the last, $q$th slot of tournament $(T,Q)$ is the European performance spot.
The allocation rule is simplified if the order of Articles 3.04 and 3.07 in the UEFA Champions League regulation (UEFA, 2024b, ) is reversed. Then the two European performance spots are allocated first (UEFA, 2024b, , Article 3.07), followed by addressing the possible vacancies created by the titleholders (UEFA, 2024b, , Article 3.04), which leads to:
$\mathcal{A}^{\prime}(\mathbf{W})=\begin{cases}i:i=p\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\geq q+1\\ i:i=r\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\leq q.\end{cases}$
Finally, any vacancy created by a titleholder can be filled by the highest-ranked not already qualified club from its association:
$\mathcal{A}^{\prime\prime}(\mathbf{W})=\begin{cases}i:i=p\lor Q_{i}(T)\leq q&\text{if }Q_{p}(T)\geq q+1\\ i:Q_{i}(T)\leq q+1&\text{if }Q_{p}(T)\leq q.\end{cases}$
Now the incentive (in)compatibility of these allocation rules can be explored.
Definition 1.
Tournament $(S,P)$ is called manipulable by team $1\leq i\leq\lvert S\rvert$ if there exists a tournament $\left(S^{\prime},P\right)$ such that $S^{\prime}=S$ except for a worse performance by team $i$, and the following two conditions hold:
  • $Q_{i}\left(S\right)=Q_{i}\left(S^{\prime}\right)$;
  • $Q_{j}\left(S\right)\geq 2$ but $Q_{j}\left(S^{\prime}\right)=1$ for a team $1\leq j\leq\lvert S\rvert$.
That is, a tanking strategy by team $i$ guarantees the same final ranking for team $i$ and makes team $j$ the tournament winner.
Definition 2.
Tournament $(T,Q)$ is called manipulable by team $1\leq i\leq\lvert T\rvert$ if there exists a tournament $\left(T^{\prime},Q\right)$ such that $T^{\prime}=T$ except for a worse performance by team $i$, and the following two conditions hold:
  • $Q_{i}\left(T\right)=Q_{i}\left(T^{\prime}\right)=q+1$;
  • $Q_{j}\left(T\right)=Q_{k}\left(T^{\prime}\right)=q-1$, as well as $Q_{j}\left(T^{\prime}\right)=Q_{k}\left(T\right)=q$ for some teams $1\leq j,k\leq\lvert T\rvert$.
That is, a tanking strategy by team $i$ guarantees that it remains ranked $(q+1)$th, while reverses teams $j$ and $k$ in the two positions directly above.
In tournament $(U,R)$, no tanking strategy exists because the UEFA club coefficients cannot be increased by a worse performance on the field.
Proposition 1.
Tournament system $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ with the allocation rule $\mathcal{A}(\mathbf{W})$ is incentive incompatible if and only if one of the following conditions hold:
  • tournament $(S,P)$ is manipulable by the team $1\leq r\leq\lvert U\rvert$ (for which $R_{r}(U)=1$) such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)\leq q-1$; or
  • tournament $(S,P)$ is manipulable by the team $1\leq i\leq\lvert T\rvert$ for which $Q_{i}(T)=q+1$ such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)=q$; or
  • tournament $(T,Q)$ is manipulable by the team $1\leq i\leq\lvert T\rvert$ for which $Q_{i}(T)=q+1$ such that the team $j$ in Definition 2 plays in both tournaments $(S,P)$ and $(T,Q)$ with $P_{j}(S)=1$, that is, $j=p$.
Proof.
If the first condition is satisfied, team $r$ can create a vacancy by making team $j$ the winner of tournament $(S,P)$, which is filled by team $r$ since $Q_{j}(T)\leq q-1$.
If the second condition is satisfied, team $i$ can create a vacancy by making team $j$ the winner of tournament $(S,P)$, which is filled by team $i$ since $Q_{j}(T)=q$.
If the third condition is satisfied, team $i$ can create a vacancy by achieving that the winner of tournament $(S,P)$ is ranked $q$th in tournament $(T,Q)$, which is filled by team $i$.
The definition of the allocation rule $\mathcal{A}$ implies that no other team can receive the possible vacant slot, thus, incentive incompatibility can arise only in the cases above. ∎
The example presented in Section 3 exploits the third condition.
Proposition 2.
Tournament system $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ with the allocation rule $\mathcal{A}^{\prime}(\mathbf{W})$ is incentive incompatible if and only if tournament $(S,P)$ is manipulable by the team $1\leq r\leq\lvert U\rvert$ (for which $R_{r}(U)=1$) such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)\leq q$.
Proof.
It follows from the proof of Proposition 1. The second and third conditions in Proposition 1 are eliminated because all cases $Q_{p}(T)\leq q$ are treated uniformly by the allocation rule $\mathcal{A}^{\prime}$. ∎
Proposition 3.
Tournament system $\mathbf{W}=\left[(S,P),(T,Q),(U,R)\right]$ with the allocation rule $\mathcal{A}^{\prime\prime}(\mathbf{W})$ is incentive incompatible if and only if tournament $(S,P)$ is manipulable by the team $1\leq i\leq\lvert T\rvert$ for which $Q_{i}(T)=q+1$ such that the team $j$ in Definition 1 plays in both tournaments $(S,P)$ and $(T,Q)$ with $Q_{j}(T)=q$.
Proof.
It follows from the proof of Proposition 1. The first and third conditions in Proposition 1 are eliminated because all cases $Q_{p}(T)\leq q$ are treated uniformly by the allocation rule $\mathcal{A}^{\prime\prime}$. ∎
The results above cannot be derived from the model of Dagaev and Sonin, (2018) since the latter contains multiple round-robin and knockout tournaments. However, if $(S,P)$ is a knockout tournament, then it cannot be manipulated by any team $1\leq i\leq\lvert S\rvert$, and both allocation rules $\mathcal{A}^{\prime}$ and $\mathcal{A}^{\prime\prime}$ guarantee incentive compatibility.
5 Discussion
@@
A recent survey on fraud in sport (Vanwersch et al.,, 2022) concludes that future research is needed to accumulate more knowledge on the prevention of fraud. Our paper examines the new format of the UEFA Champions League from this perspective. The rules have clearly improved from the 2024/25 season. Now the clubs are seeded based on their individual club coefficients except for the guaranteed place of the titleholder in the first pot (UEFA, 2024b, , Article 16.01). Therefore, the regime used until the 2014/15 season has returned, and the seeding system satisfies incentive compatibility.
@@
On the other hand, strategy-proofness is violated in the new format according to Section 3. The underlying reason is highlighted by the analysis of the mathematical model in Section 4, in particular, by Proposition 1: the European performance spots are allocated only after the vacancies caused by the two titleholders are solved (UEFA, 2024b, , Article 3.07). Thus, a team playing in the same round-robin tournament (domestic league) as one of the (potential) titleholders is interested in not creating vacancies in the UEFA Champions League. This can be achieved by “helping” other teams to be ranked over the (potential) titleholder. Such misaligned incentives may arise in the associations entitled to at least one spot in the league phase of the UEFA Champions League if they also earn one of the European performance spots. These are currently the 10 highest-ranked associations: in the 2024/25 season, England, Spain, Germany, Italy, France, the Netherlands, Portugal, Belgium, Scotland, and Austria.
Based on Section 4, incentive incompatibility is mitigated by changing the allocation rule according to the definition of either $\mathcal{A}^{\prime}$ or $\mathcal{A}^{\prime\prime}$:
@@
If rule change A is chosen, Proposition 2 remains valid. Hence, a tanking strategy can exist in the league phase of the UEFA Champions League for a team that qualifies for the champions path in the qualification of the next Champions League. For example, any team that is already eliminated in the league phase but can have the highest UEFA club coefficient in the champions path is interested in creating a vacant slot. This might be achieved by losing against a team that is expected to obtain a direct qualifying slot in its domestic league in order to help it win the Champions League. Consider the Swiss champion Young Boys in the 2024/25 UEFA Champions League. The club was eliminated in the league phase after six rounds and played against Celtic (Scotland) and Crvena Zvezda (Serbia) in the last two rounds. Since the Scottish champion directly qualifies for the league phase in contrast to the Serbian champion, Young Boys could have maximised its chances by losing against Celtic and exerting full effort against Crvena Zvezda.
On the other hand, if rule change B is chosen, Proposition 3 should be considered. Hence, a tanking strategy can exist in the league phase of the UEFA Champions League for a team if the Champions League is won by a team from the same national association due to the tanking. However, this is impossible because two teams from the same country are not allowed to play against each other in the league phase (UEFA, 2024b, , Article 16.02), and a deliberate loss in the knockout stage cannot be beneficial.
Consequently, even though rule change A seems to be closer to the original intention of UEFA, it might still lead to incentive incompatibility, albeit with a lower probability. At least, the situation presented in Section 3 will certainly be avoided. In contrast, rule change B entirely avoids misaligned incentives and highlights the role of the association constraint in the league phase draw, which is desirable not only to maximise the number of international matches played but also to ensure strategy-proofness in the league phase. This is a good illustration of the idea raised by Csató, (2022) that draw restrictions can be used to prohibit potentially unfair matches. Nonetheless, rule change B favours the domestic league of the titleholder, hence, it leads to a somewhat higher concentration of the clubs playing in the UEFA Champions League by their associations because the highest-ranked countries are more likely to win both the UEFA Champions League and the UEFA Europa League.
To summarise, the main message of Dagaev and Sonin, (2018) also applies to the model of Section 4: the vacant slots should be filled via the round-robin tournament as done by the allocation rule $\mathcal{A}^{\prime\prime}$. Therefore, UEFA could have guaranteed incentive compatibility by following the recommendation of Dagaev and Sonin, (2018) despite the different settings.
@@
If UEFA is not worried much about manipulation in a domestic league and is not willing to change its regulations, the national associations are also able to mitigate the problem to some extent. The current paper has considered incentive incompatibility as a binary concept. However, the probability of a situation that is vulnerable to tanking can be quantified via Monte Carlo situations, analogously to Csató, (2022), even if it would be difficult to implement.
Naturally, the set of matches to be played cannot be changed in a domestic championship without changing the tournament format, but the order of the matches is a decision variable. The scenario outlined in Section 3 has become threatening especially since the match Eintracht Frankfurt vs. RB Leipzig was played in the last round. The set of top teams may be relatively accurately estimated at the beginning of the season, thus, the last rounds can be scheduled to avoid any matches between the strongest teams. Such a schedule might lead to a tanking opportunity at the end of the season with a (much) lower probability. But even if Eintracht Frankfurt had already played against RB Leipzig, it would be unfair if it turns out ex post that a team has missed qualifying for the UEFA Champions League due to winning a match months before. Therefore, it is the responsibility of UEFA to act if it is really committed to its announced principles.